Evaluate the integrals.

Íx” dx = ſx=3 dx = (x-3) secx tan x dx = j(x?- Var)dx=
Answer #1


212 da- formula used canda = ht tc. Here n=12 1x12 da 12 +1 7 (213) 2 - 43 ( 27 - (0) - to ( 8112+1] 8192 tl x12da= 630-23 6x-513 da = -15 Taking constant out Scof(a) da= cffradda 2 Let u=x-3 du-da n=0 x=2 u=-3 =1 21 & du. 25 is du d 43 u-2 El da =I secx Taneda Adjusting integral based on undefined points There is an Undefined point at I If Enist b, albac, f(b) = undefinSecca) Tan (n) da=diverges ř G2_ſa) da. Berda- [ sudx. serde-I = 2 de 312 312 - (+3) » – ₃ (+212) 5(8-0) - ( 2312-0) - 0.7810